Write an equation for the line in point/slope form and slope/intercept form that has the given condition.

4. Slope= 3/2 and passes through the origin

5. X-intercept=4 and Y-intercept=-3
1
Expert's answer
2013-01-29T11:22:49-0500
4. general form for equation of a straight line is y=mx+bif the slope is (3/2) then m=3/2, ifthe line passes through origin b=0.so we obtain y=(3/2)x
5. line goes through (4, 0), (0, -3);
(x-4)/-4=y/3 => 3(x-4)=-4y => 3x+4y=12;